Operations Report – April 13th

Crew 245 Operations Report 13-04-2021

SOL: 3

Name of person filing report: Shravan Hariharan

Non-nominal systems: Nothing to report

Notes on non-nominal systems: Nothing to report

Generator: Not run, as solar power was sufficient

Hours run: N/A

From what time last night: N/A

To what time this morning: N/A

List any additional daytime hours when the generator was run: N/A

Solar— SOC 80% (Before generator is run at night)

Diesel Reading – 50%

Station Propane Reading – ?% – Director, please confirm

Water (loft tank): 25 gallons

Water Meter: 150273.7 units

Water (static tank): ? gallons

Static to Loft Pump used – yes

Water in GreenHab: 0 gallons

Water in ScienceDome: 0 gallons

Toilet tank emptied: no

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: N/A

Spirit rover used: no

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: yes, in town

Opportunity rover used: no

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: no, in town

Curiosity rover used: no

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: yes, in town

Perseverance rover used: Yes

Hours: 200.3

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Notes on rovers: Nothing to report

ATV’s Used: 350.1

Reason for use: EVA

Oil Added? No

ATV Fuel Used: N/A

# Hours the ATVs were used today: 2.5 (but 0.5 hours of active transit)

Notes on ATVs: Mission Director will monitor and refuel ATVs for duration of mission

HabCar used and why, where?: No

CrewCar used and why, where?: No

General notes and comments: Hab kitchen sink is leaking, and seal will be re-caulked around border of sink. Leak is minor, and should not require any outside assistance.

Summary of internet: Crew has almost entirely consumed data for the day, so additional monitoring and rationing needed.

Summary of suits and radios: All radios nominal, suits currently charging. Suit batteries successfully swapped by Crew Engineer last night, and all suits functioned properly on EVAs today.

Summary of Hab operations: Nothing to report

Summary of GreenHab operations: None

Summary of ScienceDome operations: Used by Crew Botanist and Crew Scientist for botany experiments

Summary of RAM operations: Used by Crew Scientist to build hydroponics system

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: HSO report coming in with minor notes

Questions, concerns and requests to Mission Support: How do we check the amount of water in the static tank? We were unable to find a meter.


Shravan Hariharan, Crew Engineer
Red Planet People – MDRS Crew 245 "Team Patamars"
To Mars and Beyond – For All!